I need help calculating Forward rate and interest rate parity for INR/USD

How to calculate the nominal interest rate, I$ and I rupee?

where to obtain the spot exchange rates for USD and INR

How to calculate 1 forward rate
Answer #1
(a) Forward Rate(f/d) = Spot Rate(f/d) X [(1 + if) / (1 + id)]

f is foreign currency, d is domestic currency, if is interest rates in foreign currency and id is interest rates in domestic currency

Keep 'f' as INR and 'd' as USD in above equation will give the forward rates given spot rates and interest rates in both countries

(b) Nominal interest rate doesn't consider inflation into account but real interest rate does, nominal means only in monetary terms without effect of inflation.

The exact formula is (1+real interest rate) * (1+ inflation rate) = (1+nominal interest rate)

The close formula is Real interest rate + inflation rate = Nominal interest rate

Please note that real interest rates do not change so if inflation increases, nominal interest rates will also increase.

Again for Dollar and Rupee, real interest rates in both currencies will remains same, so re align the above equation gives below formula

( 1+ nominal interest rate (rupee))/(1+inflation (rupee)) = ( 1+ nominal interest rate (dollar))/(1+inflation (dollar))

c) For spot exchange rates, you can always contact banks or some external reliable websites

d) (1 + 1 year spot rate) * (1+ 1 year forward rate) = ( 1 + 2 year spot rate)^2

This equation says if I invest my money for two years, but in two parts. One year from now and one more year when my first year ends is equal to if i invest my money for two straight years. The final amount will be same in both cases. This can be used for any number of years to calculate respective forward rate.